A random sample of 75 measurements is drawn from a population with a mean of 60 and a standard deviation of 5.

17.) The mean of the sampling distribution of y hat is ________.

18.) The standard error of the sampling distribution of y hat is __________. (use 4 decimal places).
Answer #1

Question 17

= µ = µ = 60

Question 18

σ = σ / √ (n) = 5/√75 = 0.5774